A 100 foot wire is extended from the ground to the top of a 60 foot pole, whish is perpendicular to the level ground. Find to the nearest degree, the measure of the angle that the wire makes with the ground.


1
Expert's answer
2021-01-07T13:47:43-0500


Here in the image, AB is the 60ft tall pole and AC is the 100ft tall wire.

As "\angle" ABC is "90^o" , we use trigonometry:


"sin\theta=\frac{AB}{AC}=\frac{60ft}{100ft}=\frac{60}{100}=\frac{3}{5}"

Therefore we have,

"sin\theta=\frac{3}{5}\\
\Rightarrow \theta=sin^{-1}(\frac{3}{5})\\
\Rightarrow \theta\approx0.6435\approx36.8698^o"




Therefore the angle that the wire makes with the ground is "36.8698^o."